OpenStudy (anonymous):

IS THIS RIGHT? Find the area bounded by y = cosx, y = sinx and the y axis solution below

OpenStudy (anonymous):

\[A = cosx dx- sinx dx A= sinx - (-cosx) A = \sin(\pi/4) - \sin(0) - ( cospi/4-\cos0) A = 1/\sqrt{2} - 0[- 1/\sqrt{2} - 1]\]

OpenStudy (anonymous):



OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@nikvist

OpenStudy (anonymous):

and how to graph that one @nikvist

OpenStudy (nikvist):

\[S=\int\limits_{0}^{\pi/4}\left(\cos{x}-\sin{x}\right)dx\]

OpenStudy (anonymous):

how o graph that y - axis, cosx and sinx

OpenStudy (anonymous):

and i have question why the upper limit become pi/4 and the lower limit become 0?

OpenStudy (anonymous):

@nikvist

OpenStudy (anonymous):

why 0 become the lower limit?

OpenStudy (anonymous):

is that the graph of all?

OpenStudy (anonymous):

@nikvist

OpenStudy (nikvist):

x=0 (y-axis) is a begin of this area.

OpenStudy (anonymous):

why pi/4 become pi/4?

OpenStudy (nikvist):

Because x=pi/4 is end of area. Intersection of graph sin(x) and cos(x) is point (pi/4,sqrt(2)).

OpenStudy (nikvist):

(pi/4,sqrt(2)/2)

OpenStudy (anonymous):

how about the 0?

OpenStudy (anonymous):

cant understand it easily

OpenStudy (anonymous):

im havig difficulties in figuring out how pi/4 and 0 put in upper and lower limit

OpenStudy (nikvist):

Simple, this area bounded: left x=0, right x=pi/4, up y=cos(x), down y=sin(x). See graph.

OpenStudy (anonymous):

the graph is incomplete , theres no y axis graph

OpenStudy (mathmate):

 Above graph shows what @nikvist had explained.